No of real solution of  equation (15+4*14^0.5)^t   +(15-4*14^0.5)^t  = 30  where t= x^2 – 2*|x| is …..

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Dear Student,
let (15+4*14^0.5)^t = a
         Then the given equation is a + 1/a = 30
         Solving for a and then t we get t=1
         X^2 – 2|x| = 1
  • |x|= 1+√5
  • X = -1-√5 , 1+√5
2 solutions 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)